# Adaptive Grid-Based Thompson Sampling for Efficient Trajectory Discovery

### v1

@@ 5.2 Calibrating to both hospitalizations and deaths
As in previous work, we compare model outputs to the two empirical datasets that were the most reliably reported and available during the COVID-19 pandemic: the daily count of hospitalizations and deaths attributed to COVID-19 from April through June 2020, which was accessed through the Chicago Data Portal (CityOfChicago2021, ). Let $h_{t}$ and $d_{t}$ be the empirical counts of deaths and hospitalizations, respectively, at time $t$. Since we are concerned not just with parameters $\bm{\theta}$ but also the random seed $r$, let $\bm{\theta}_{r}:=(\theta_{1},\theta_{2},\theta_{3},r)$ specify inputs that generate a particular CityCOVID trajectory. Let $h_{\bm{\theta}_{r},t}$ and $d_{\bm{\theta}_{r},t}$ denote the simulated count of hospitalizations and deaths. We define the dual-objective function
@@ 5.4 Results
Further examination of the regions of parameter space that are explored for the $\mathcal{L}(\theta_{r})<20$ threshold shows that aCRN finds trajectories across a much more diverse region, capturing multiple plausible solutions, whereas fHet concentrates almost exclusively on a single grid point in the fixed grid, repeatedly sampling additional replicates around it. This highlights the advantage of combining the CRNGP surrogate with adaptive-grid refinement, allowing aCRN to efficiently balance exploration of new parameter regions with exploitation of high-likelihood areas (see Figure 23 in the Supplementary Materials).
@@ 6 Conclusion and Future Work
Several avenues for future work remain. First, strategies for selecting and updating the replicate set could be extended beyond the fixed design considered here. Systematic growth of the replicate set, guided by updated GP draws (chevalier2015fast, ) or their continuous approximations (mutny2018efficient, ), would increase flexibility while maintaining computational tractability. Second, explicit model-discrepancy terms could be integrated into the optimization framework to account for differences between simulator outputs and observed data (koermer2024augmenting, ). Third, the current CRNGP covariance function assumes a uniform similarity index across seeds (Equation 6); more general formulations could leverage grouping structures for categorical inputs (roustant2020group, ) or exploit more complex relationships between seeds in latent spaces (da2025distributional, ). Finally, while our framework focused on scalar summary outputs, extensions to functional or temporal outputs may yield further improvements, as demonstrated in related work on multi-output and functional BO (astudillo2021thinking, ). Although motivated here by epidemic modeling, the proposed framework is general and directly applicable to a wide range of fields that employ stochastic simulators, where realization-level calibration may be beneficial for accurate inference and decision support.

### v2

@@ 5.2 Calibrating to Both Hospitalizations and Deaths
As in previous work, we compare model outputs to the two empirical datasets that were the most reliably reported and available during the COVID-19 pandemic: the daily count of hospitalizations and deaths attributed to COVID-19 from April through June 2020, which was accessed through the Chicago Data Portal City of Chicago (2021). Let $h_{t}$ and $d_{t}$ be the empirical counts of deaths and hospitalizations, respectively, at time $t$. Since we are concerned not just with parameters $\bm{\theta}$ but also the random seed $r$, let $\bm{\theta}_{r}:=(\theta_{1},\theta_{2},\theta_{3},r)$ specify inputs that generate a particular CityCOVID trajectory. Let $h_{\bm{\theta}_{r},t}$ and $d_{\bm{\theta}_{r},t}$ denote the simulated count of hospitalizations and deaths. We define the dual-objective function
@@ 5.4 Results
Further examination of the regions of parameter space that are explored for the $\mathcal{L}(\theta_{r})<15$ threshold shows that aCRN finds trajectories across a much more diverse region, capturing multiple plausible solutions, whereas fHet concentrates almost exclusively on a single grid point in the fixed grid, repeatedly sampling additional replicates around it. This highlights the advantage of combining the CRNGP surrogate with adaptive-grid refinement, allowing aCRN to efficiently balance exploration of new parameter regions with exploitation of high-likelihood areas (see Figure S16 in the Supplementary Materials).
@@
Figure 14: Parameter-replicate tuples corresponding to the trajectories with $\mathcal{L}(\bm{\theta}_{r})<15$ found by aCRN and fHet.
@@ 6 Conclusion and Future Work
Several avenues for future work remain. First, strategies for selecting and updating the replicate set could be extended beyond the fixed design considered here. Systematic growth of the replicate set, guided by updated GP draws Chevalier et al. (2015) or their continuous approximations Mutny and Krause (2018), would increase flexibility while maintaining computational tractability. Second, explicit model-discrepancy terms could be integrated into the optimization framework to account for differences between simulator outputs and observed data Koermer et al. (2024). Third, the current CRNGP covariance function assumes a uniform similarity index across seeds (Equation 6); more general formulations could leverage grouping structures for categorical inputs Roustant et al. (2020) or exploit more complex relationships between seeds in latent spaces Da Veiga (2025). Finally, while our framework focused on scalar summary outputs, extensions to functional or temporal outputs may yield further improvements, as demonstrated in related work on multi-output and functional BO Astudillo and Frazier (2021). Although motivated here by epidemic modeling, the proposed framework is general and directly applicable to a wide range of fields that employ stochastic simulators, where realization-level calibration may be beneficial for accurate inference and decision support.